OpenStudy (anonymous):

Econ-math question. "lf we set the absolute value of the MRS(marginal rate of substitution) equal to the price ratio, we have the equation..." So my mrs is -(y+1)/x+2 and my price ratio is simply one. From there the instructor wrote (y+1)/(x+2) = 1, which makes sense because absolute value. He then has that => x-y=-1 and that's where I'm confused.

OpenStudy (phi):

\[ \frac{(y+1)}{(x+2)} = 1 \] multiply both sides by (x+2) \[ \frac{(y+1)}{(x+2)}\cdot (x+2) = x+2 \] and simplify

OpenStudy (phi):

y+1 = x+2 add -y to both sides 1 = x - y + 2 add -2 to both sides -1 = x -y

OpenStudy (anonymous):

oy vey, i dont know why that was so hard for me to wrap my head around. thank you.

OpenStudy (phi):

You won't get tricked by that one anymore.

OpenStudy (anonymous):

Absolutely won't.